zarasht:

what is the mass in grams of KBr in 0.400 L of a 0.350 M solution???

dude:

Okay, so first you want to know the moles in the KBr solution - They told you this is a 0.350 M solution \(.400\color{red}{\cancel{L}}\times\dfrac{0.350M}{\color{red}{1\cancel{L}}}=\) From then you need to multiply what you get from above by the molar mass (Which is 119.002) \(Moles\times Molar~mass \rightarrow Moles \times 119.002=\)

zarasht:

I got 16. 7 as my answer. Would that be right?

dude:

Yes 16.7 grams (rounded)

zarasht:

okay thank you